The saturation current will be the twice of the diode current. and contact potential of the diode is 33 Volt. calculate the temperature.


1
Expert's answer
2021-06-15T00:57:01-0400

Ideal diode equation is given by

I=I_o(e^{\frac {qV}{kT}}-1)

Diode current=I

Saturation current= I_o

Electron charge=q

Contact voltage=V

Boltzmann's constant=k

Temperature= T

Making T the subject of the formula

T= \frac {\frac {qV}{k}}{\ln(\frac{I}{I_o}+1)}

q=1.6 \times 10^{-19}C

V=33v

k= 1.38 \times 10^{-23}m^2kgs^{-2}K^{-1}

I_o=2\times I

T= \frac {\frac {1.6 \times 10^{-19} \times 33}{1.38 \times 10^{-23}}}{\ln(\frac{I}{2 \times I}+1)}

T= \frac{382608.69}{\ln1.5}

T=943629K